29. Write a rule for the function:x=2,3,7,10 y=8,7,3,0
30. Write a rule for the function: x=0,4,8,12 y=5,7,9,11

Answer by stanbon(75887)

29. Write a rule for the function:x=2,3,7,10 y=8,7,3,0
Notice that when x goes down 1, the y value goes down one.
When x goes down 4, the y value goes down four.
etc.
Therefore the function is linear and looks like y = mx+b
So take two of the points and find the equation:
Take (2,8) and (10,0)
m = (8-0)/(2-10) = -1
Since m=-1 and y=8 when x = 2 you can find "b", as follows:
8 = -1*2+b
b=10
-----------
EQUATION:
y = -x+10
===========================
30. Write a rule for the function: x=0,4,8,12 y=5,7,9,11
Try to see the change pattern.
As x increases 4, y decreases 2
The function must be linear: y=mx+b
Pick two points; find the slope; find b.
Write the equation:
y = (-1/2)x + 5
========================
Cheers,
Stan H.